# Blackbox Dataset Inference for LLM

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 32nd ACM Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Large language models (LLMs), such as the Generative Pre-trained Transformer (GPT) (floridi2020gpt, ), have made remarkable progress in natural language processing over recent years. LLMs are trained on vast amounts of text data, enabling them to generate human-like text, answer questions, and perform various language-related tasks. LLMs are initially trained on a broad dataset and then fine-tuned for specific tasks. This approach has proven highly effective in enhancing performance across a wide range of language tasks (min2023recent, ).
@@
Problem: Recently, the training of LLMs using massive public data has sparked significant privacy concerns. The inclusion of personally identifiable information and copyrighted material in training datasets has led to legal conflicts, such as the lawsuit between The New York Times and OpenAI (lawsuit, ) and the suspension of ByteDance’s GPT accounts (bytedanceopenai, ). These disputes underscore the problem of using copyrighted content without proper attribution or licensing, potentially infringing on the rights of the data creators. For simplicity, we define this problem as dataset misuse.
@@
Literature: To mitigate the problem of dataset misuse, there are two applicable solutions. One approach is membership inference attacks (MIAs) (shokri2017membership, ). MIA aims to determine if a specific data sample is a part of the training dataset. Applying LLM-level MIA (shi2023detecting, ; mattern2023membership, ) to all samples in a suspect dataset $\mathcal{D}$, we can validate how frequently the samples are used to train the target model and, thus, infer dataset misuse. However, previous research (maini2021dataset, ) has highlighted an impossibility result, indicating that as the training dataset size increases, the accuracy of membership inference diminishes to the level of random guessing. Further, a recent study (maini2024llm, ) unveils that MIAs against LLMs present random accuracy when the training set and testing set are independent and identically distributed (IID).
@@
Another method is dataset inference (maini2024llm, ). The idea is to aggregate the results of multiple MIAs across different data samples to detect if a suspect model $\mathcal{M}$ used a victim dataset $\mathcal{D}$ in training. Briefly, the method builds a set of diverse MIAs and trains a linear regression model to learn which MIAs are more helpful. Applying the linear model and statistical t-test, the method determines if $\mathcal{M}$ can better “recognize” $\mathcal{D}$ than a validation dataset that is IID with $\mathcal{D}$. For successful dataset inference, the method requires MIAs to present a certain level of accuracy and, thus, assumes grey-box access to $\mathcal{M}$ for obtaining rich inputs (probabilities, loss, perplexity, etc.) for MIAs. This brings limitations on practicality. LLMs, especially those deployed for profits, have limited incentives to return the intermediate outputs. For instance, Google’s Gemini (googlemodels, ) and Anthropic’s Claude (Anthropicmodels, ) only offer text-based responses when queried with web-based interfaces or APIs.
@@
Our Method: In this paper, we focus on black-box dataset inference for LLMs. In other words, we aim to determine if a suspect model $\mathcal{M}$ used a victim dataset $\mathcal{D}$ in training, entirely relying on the text-based responses of $\mathcal{M}$. Our key observation is that certain data samples, when involved in training, drive the model to generate responses highly resembling the oracle output (i.e., the output coming with the samples). We define those samples as tainted samples, and we develop a three-step method exploiting tainted samples for dataset inference. ❶ We build a set of reference models that are certainly not trained on $\mathcal{D}$, annotated as $\mathbf{R}$ = $\{\mathcal{R}_{1},\mathcal{R}_{2},...,\mathcal{R}_{n}\}$. We fine-tune each reference model with $\mathcal{D}$ to produce trained reference models $\mathbf{R}^{t}$ = $\{\mathcal{R}_{1}^{t},\mathcal{R}_{2}^{t},...,\mathcal{R}_{n}^{t}\}$. ❷ We pre-process $\mathcal{D}$ to identify candidates of tainted samples based on the similarity between the oracle outputs and the responses from $\mathbf{R}$ and $\mathbf{R}^{t}$. ❸ We compare $\mathcal{M}$ with $\mathbf{R}$ and $\mathbf{R}^{t}$ on tainted samples, determining if $\mathcal{M}$ used $\mathcal{D}$ according to which reference set $\mathcal{M}$ is closer to.
@@
Evaluation: We perform a series of evaluations, involving member and non-member models collected from the wild and built locally, to assess the performance of our method. They show that our method can achieve 100% accuracy under the non-IID setting, given different target datasets. Our method also presents very high accuracy under the IID setting, missing only a few cases. By further increasing the reference models and diversifying their architectures, we can amend our method to correct those cases. In contrast, the baseline based on blackbox MIA (dpdllm, ) fails to offer satisfactory accuracy under either non-IID or IID settings. Oftentimes, the baseline even approximates a random guess.
We further experiment with various evasion attempts against our method, including rephrasing responses, varying temperature, removing tainted samples, and training on subsets. The results show that our method preserves its accuracy under those attempts. The only exception occurs when the adversaries only train their models using fewer than 50% of the samples, which we argue is no longer a problem suitable to be solved by dataset inference (see §6.4). These provide empirical evidence for the robustness of our method.
@@
  • We propose a new method to detect dataset misuse in LLM training. Compared to the previous methods, our method only needs black-box access to the target model.
@@
  • We run intensive evaluations of our method. The results show that our method offers high accuracy in all settings and presents robustness against evasion attempts.
2. Background
2.1. Large Language Models
In the realm of artificial intelligence, large language models (LLMs) have emerged as transformative tools capable of understanding and generating human-like texts. LLMs typically take in a prompt, such as “write a poem”, and produce a text that satisfies the requirement, i.e., a poem. More technically, LLMs encode and pass the prompt through several embedding layers to generate a probability distribution over the possible next tokens. The token with the highest probability or based on other strategies (greedy_beam, ; topk, ; topp, ) is chosen as the prediction. The predicted token is added to the sequence, and the model repeats the process to generate subsequent tokens.
Today’s LLMs share a hybrid ecosystem regarding openness. A myriad of LLMs, including Meta’s Llama family (llama, ), Google’s Gemma family (team2024gemma, ), and the Mistral family (jiang2023mistral, ), release their source code and models for free. Many of them, such as Mistral 7B, even adopt permissive licenses like Apache 2.0 to allow unrestricted commercialization (openllms, ). Yet, many LLMs, with OpenAI’s GPT series (openaimodels, ), Google’s Gemini series (googlemodels, ), and Anthropic’s Claude series (Anthropicmodels, ) as examples, are developed for commercial use. They only open web- or API-based interfaces to support prompt-response interactions.
2.2. LLM Training
Training an LLM typically starts with pre-training on a diverse corpus of text data to build a base model. This base model generally understands natural language but is not optimized for executing tasks. Thus, it can generate contextually appropriate responses to the prompt but may not strictly adhere to specific instructions.
A base model can be optimized through fine-tuning on a dataset consisting of instructions and responses to follow those instructions (howard2018universal, ; devlin2018bert, ; radford2018improving, ). The fine-tuning process derives an instruct model, which can interpret instructions accurately and act accordingly. For illustration, we show a comparison between a base model and an instruct model in Figure 1. An instruct model can be further fine-tuned for optimizations. For example, OpenAI offers APIs for fine-tuning various GPT-based instruct models (GPT-3.5-Turbo, GPT-4, GPT-4o, etc.) (finetuneapi, ).
Prompt: What is the capital of France? Base Model (DaVinci-002): “A. Paris B. France C. Lyon D. Rouen.” Instruct Model (GPT-3.5-Turbo): “Paris”
Figure 1. Base model v.s. instruct model
2.3. Dataset Misuse
The availability of powerful hardware enables broad development of LLMs in various domains. Fine-tuning, allowing anyone to customize LLMs at a low cost, intensifies this trend. Together with the advancement comes ethical or legal concerns about the training data. For instance, many datasets on the market (e.g., Alpaca  (Alpaca_dataset, ) and SlimOrca (Slimorca_dataset, )) are composed of outputs by OpenAI’s GPT models, which have been frequently used to train LLMs (DeciFoundationModels, ; lian2023jackalope, ; hao2024exploring, ; model_007_13b_v2, ). When deployed for commercial purposes, these LLMs compromise OpenAI’s terms of use, which explicitly disallow “using our output to develop models that compete with OpenAI” (Termsofu22:online, ). On this ground, OpenAI suspends ByteDance’s account after it allegedly used GPT to build rival AI products in 2023 (bytedanceopenai, ).
In general, the problem is misuse of publicly accessible but copyright-protected datasets for training LLMs, especially when the LLMs are trained for profits. To mitigate dataset misuse, a promising direction is to develop methods to detect if an LLM’s training involves a victim dataset (maini2024llm, ).
3. Threat Model
@@
In this paper, we focus on detecting dataset misuse in LLM training. Following the settings of recent model ownership verification (zhou2024revisiting, ; grove, ), we assume a scheme with three participants involved, including $Accuser$, $Arbiter$, and $Responder$. $Accuser$ owns a publicly-accessible dataset $\mathcal{D}$ and disallows using $\mathcal{D}$ to train LLMs for profits. $Responder$ owns a private LLM $\mathcal{M}$ and allows paid queries to $\mathcal{M}$. Believing that $Responder$ secretly trained or fine-tuned $\mathcal{M}$ on $\mathcal{D}$, $Accuser$ makes an accusation of dataset misuse. $Arbiter$, a trusted third party, aims to verify the accusation.
@@
4. Motivation and Key Observations
4.1. Previous Research
Maini et. al. generalize the problem of detecting dataset misuse by LLMs as dataset inference (maini2024llm, ). For consistent references, we will reuse this concept in this paper. Notably, the same concept has also been used in the context of model ownership resolution (maini2021dataset, ), which is out of our scope.
@@ 4.1. Previous Research
Background: To achieve dataset inference, Maini et al. propose the first yet only method (maini2024llm, ), which we name as DIAα for easy references. DIAα is built on top of MIAs (shokri2017membership, ) against LLMs (shi2023detecting, ; mattern2023membership, ). At the high level, MIA is a function $f_{\mathcal{M}}:\mathcal{X}\rightarrow\{0,1\}$. Given an LLM model $\mathcal{M}$ and an input $x$ from space $\mathcal{X}$, $f_{\mathcal{M}}$ determines if $x$ is in $\mathcal{M}$’s training set (i.e., member or not). A representative $f_{\mathcal{M}}$ is threshold-based: $f_{\mathcal{M}}=\mathbb{1}[\mathcal{S}(\mathcal{M},x)<\delta]$, where $\mathcal{S}$ is a score function like loss (yeom2018privacy, ) and perplexity (carlini2021extracting, ) and $\delta$ is a pre-determined threshold. There are also many other variants of $f_{\mathcal{M}}$, including lowest probability-based (shi2023detecting, ), perturbation based (maini2021dataset, ; mitchell2023detectgpt, ), reference model based (maini2024llm, ), and entropy based (carlini2021extracting, ).
@@
DIAα Design: Observing that MIAs may present random accuracy when detecting membership of individual data samples (especially when the training set and testing set of $f_{\mathcal{M}}$ are IID, namely independent and identically distributed (ganssler1979empirical, )), DIAα aggregates the results of multiple MIAs across different data samples to perform dataset inference. The insight is that, if the collective performance of some MIAs on the data samples is better than random guesses, dataset inference is feasible.
@@
Technically, to perform inference of a victim dataset $\mathcal{D}$ against a suspect model $\mathcal{M}$, DIAα gathers a validation dataset $\hat{\mathcal{D}}$ (which is IID with $\mathcal{D}$) and builds a set of diverse MIAs. Each MIA is trained with a training split from $\mathcal{D}$ as members and a training split from $\hat{\mathcal{D}}$ as non-members. Using the outputs of all MIAs as features and the membership status as labels, DIAα further trains a linear regression model, aiming to learn which MIAs are more helpful. Another split from $\mathcal{D}$, the testing split, is then fed to the linear model to obtain a membership likelihood value for each sample, forming a likelihood vector. Similarly, a likelihood vector can be obtained with the testing split from $\hat{\mathcal{D}}$. Given the two likelihood vectors, DIAα performs t-test (kim2015t, ) using the alternate hypothesis that $\mathcal{D}$ is used for training $\mathcal{M}$, or mathematically, the mean value of $\mathcal{D}$’s likelihood vector is higher than $\hat{\mathcal{D}}$’s.
@@
DIAα Restrictions: For successful dataset inference, DIAα requires MIAs to present a certain level of accuracy. To this end, DIAα assumes grey-box access to $\mathcal{M}$ for obtaining rich inputs for MIAs. Specifically, the MIAs adopted by DIAα all require intermediate outputs from $\mathcal{M}$ (probabilities, loss, perplexity, etc.) to predict a data sample’s membership (R1). This reduces DIAα’s practicability. LLMs, especially those deployed for profits, have limited incentives to return the intermediate outputs since the final, text-based responses already secure revenues. For instance, Google’s Gemini (googlemodels, ) and Anthropic’s Claude (Anthropicmodels, ) only offer text-based responses when queried with web-based interfaces or APIs.
@@
To augment the capacity of dealing with IID datasets, DIAα requires the availability of a validation dataset that is IID with the victim dataset (R2). This increases the burden of using DIAα. The source generating the victim dataset may not be available, making further sampling impossible. Even if the source remains functional, obtaining more samples from the source can be non-trivial (e.g., getting more books from the same author is not always possible).
4.2. Our Observations
@@
In principle, MIAs exploit group-level discrepancies between the training set and the non-training set. For instance, a representative threshold-based MIA (yeom2018privacy, ) relies on a clear separation between the loss of training and non-training samples. When the group-level discrepancies do not emerge, MIAs fail. Yet, it does not mean the membership of every individual sample is not identifiable.
@@
Table 1. Tainted samples identified from databricks-dolly-15k when used to fine-tune GPT-3.5-Turbo.
@@
Figure 2. A comparison of GPT-3.5-Turbo’s response before and after fine-tuning. The longest common sub-sequence shared with the oracle output is highlighted in green.
❶ Study Setup: We fine-tune GPT-3.5-Turbo with databricks-dolly-15k (Databricks_dataset, ), a dataset of 15,011 instruction-following records published by Databricks in April 2023. The training data of GPT-3.5-Turbo was up to September 2021 (gpt3turbo, ), ensuring databricks-dolly-15k was not included. We use the official APIs offered by OpenAI (openai_fine_tuning, ) to perform fine-tuning. We set the training epoch (n_epochs) to be 3, leaving all other parameters to default.
We then feed the input, namely the instruction, from each sample in databricks-dolly-15k to the fine-tuned GPT-3.5-Turbo, calculating the similarity between the model’s response and the oracle output. We adopt BERTScore (zhang2019bertscore, ) as the similarity metric, which leverages the pre-trained contextual embeddings from BERT (devlin2018bert, ) and matches words in the response and oracle ouput by cosine similarity. BERTScore offers two advantages. First, it has been shown to correlate with human judgment on sentence-level and system-level evaluation (berthugginface, ). Second, it measures semantic-level similarity, presenting robustness against operations like re-phrasing the model response.
❷ Results: In the study, we only consider samples resulting in a BERTScore over 0.95 to be tainted samples. Even using this conservative standard, we identify hundreds of tainted samples from the major categories in databricks-dolly-15k, as summarized in Table 1. We also demonstrate an example in Figure 2, attaching the model’s responses before and after fine-tuning for references. Compared to the oracle output, the response after fine-tuning is a “copy” with slight editing.
@@
Figure 3. A databricks-dolly-15k example where GPT-3.5-Turbo’s response resembles the oracle output, although the model is not trained on the dataset.
Table 2. “Fake” tainted samples identified from databricks-dolly-15k with non-fine-tuned GPT-3.5-Turbo.
@@
Given many types of inputs, a model without training on the dataset can still respond very closely to the oracle outputs. Figure 3 presents an illustrative example, which we may mistakenly classify as a tainted sample even using a high similarity threshold. Such cases are common. We redo the study presented in Observation I but using GPT-3.5-Turbo without fine-tuning. As summarized in Table 2, hundreds of samples from databricks-dolly-15k are classified as tainted samples. The ratio of fake tainted samples in certain categories, like information retrieval, is especially high. In this case, the method above will wrongly report that GPT-3.5-Turbo used databricks-dolly-15k for training.
@@
Discussion: While we cannot reliably pinpoint individual tainted samples, we may still rely on the statistical results of tainted samples for dataset inference. As shown in Table 1 and Table 2, the total number of tainted samples after fine-tuning is significantly higher than before fine-tuning. This inspires a new idea to determine if suspect model $\mathcal{M}$ used victim dataset $\mathcal{D}$ for training: if the amount of tainted samples identified with $\mathcal{M}$ is closer to the situation where $\mathcal{D}$ is involved for training, we report $\mathcal{M}$ used $\mathcal{D}$, and not otherwise.
5. Our Methods
5.1. Overview

### v3

@@ Abstract
I Introduction
@@
Large language models (LLMs), such as the Generative Pre-trained Transformer (GPT) [1], have made remarkable progress in natural language processing over recent years. LLMs are trained on vast amounts of text data, enabling them to generate human-like text, answer questions, and perform various language-related tasks. LLMs are initially trained on a broad dataset and then fine-tuned for specific tasks. This approach has proven highly effective in enhancing performance across a wide range of language tasks [2].
@@
Problem: Recently, the training of LLMs using massive public data has sparked significant privacy concerns. The inclusion of personally identifiable information and copyrighted material in training datasets has led to legal conflicts, such as the lawsuit between The New York Times and OpenAI [3] and the suspension of ByteDance’s GPT accounts [4]. These disputes underscore the problem of using copyrighted content without proper attribution or licensing, potentially infringing on the rights of the data creators. For simplicity, we define this problem as dataset misuse.
@@
Literature: To mitigate the problem of dataset misuse, there are two applicable solutions. One approach is membership inference attacks (MIAs) [5]. MIA aims to determine if a specific data sample is a part of the training dataset. Applying LLM-level MIA [6, 7] to all samples in a suspect dataset $\mathcal{D}$, we can validate how frequently the samples are used to train the target model and, thus, infer dataset misuse. However, previous research [8] has highlighted an impossibility result, indicating that as the training dataset size increases, the accuracy of membership inference diminishes to the level of random guessing. Further, a recent study [9] unveils that MIAs against LLMs present random accuracy when the training set and testing set are independent and identically distributed (IID).
@@
Another method is dataset inference [9]. The idea is to aggregate the results of multiple MIAs across different data samples to detect if a suspect model $\mathcal{M}$ used a victim dataset $\mathcal{D}$ in training [9]. Briefly, the method builds a set of diverse MIAs and trains a linear regression model to learn which MIAs are more helpful. Applying the linear model and statistical t-test, the method determines if $\mathcal{M}$ can better “recognize” $\mathcal{D}$ than a validation dataset that is IID with $\mathcal{D}$. For successful dataset inference, the method requires MIAs to present a certain level of accuracy and, thus, assumes grey-box access to $\mathcal{M}$ for obtaining rich inputs (probabilities, loss, perplexity, etc.) for MIAs. This brings limitations on practicality. LLMs, especially those deployed for profits, have limited incentives to return the intermediate outputs. For instance, Google’s Gemini [10] and Anthropic’s Claude [11] only offer text-based responses when queried with web-based interfaces or APIs.
@@
Our Method: In this paper, we focus on black-box dataset inference for LLMs. In other words, we aim to determine if a suspect model $\mathcal{M}$ used a victim dataset $\mathcal{D}$ in fine-tuning, entirely relying on the text-based responses of $\mathcal{M}$. Our key observation is that certain data samples, when involved in fine-tuning, drive the model to generate responses highly resembling the oracle output (i.e., the output coming with the samples). We define those samples as tainted samples, and we develop a three-step method exploiting tainted samples for dataset inference. ❶ We build a set of non-member reference models that are certainly not trained on $\mathcal{D}$, annotated as $\mathbf{R}$ = $\{\mathcal{R}_{1},\mathcal{R}_{2},...,\mathcal{R}_{n}\}$. We fine-tune each reference model with $\mathcal{D}$ to produce member reference models $\mathbf{R}^{t}$ = $\{\mathcal{R}_{1}^{t},\mathcal{R}_{2}^{t},...,\mathcal{R}_{n}^{t}\}$. ❷ We pre-process $\mathcal{D}$ to identify candidates of tainted samples based on the similarity between the oracle outputs and the responses from $\mathbf{R}$ and $\mathbf{R}^{t}$. ❸ We compare $\mathcal{M}$ with $\mathbf{R}$ and $\mathbf{R}^{t}$ on tainted samples, determining if $\mathcal{M}$ used $\mathcal{D}$ according to which reference set $\mathcal{M}$ is closer to.
@@
Evaluation: We perform a series of evaluations, involving member suspect models fine-tuned with $\mathcal{D}$ and non-member suspect models not fine-tuned with $\mathcal{D}$ collected from the wild and built locally, to assess the performance of our method. They show that our method can achieve perfect accuracy under the non-IID setting, given different victim datasets. Our method also presents very high accuracy under the IID setting, missing only a few cases. By further increasing the reference models and diversifying their architectures, we can amend our method to correct those cases. In contrast, the baseline based on blackbox MIA [12] [13] fails to offer satisfactory accuracy under either non-IID or IID settings. Oftentimes, the baseline even approximates a random guess. Additionally, we evaluate our method without using tainted samples, and the results demonstrate their necessity.
We further experiment with various evasion attempts against our method, including rephrasing responses, varying temperature, removing tainted samples, training on subsets, and training on multiple datasets sequentially. These provide empirical evidence for the robustness of our method. Furthermore, additional experiments with datasets of varying sizes and formats also demonstrate the generalization capability of our method.
@@
  • We propose a new method to detect dataset misuse in LLM training. Compared to the previous methods, our method only needs black-box access to the suspect model.
@@
  • We run intensive evaluations of our method. The results show that our method offers high accuracy in all settings, presents robustness against evasion attempts, and provides generalization to various scenarios.
II Background
II-A Membership Inference
Membership inference attacks (MIAs) aim to determine whether a specific sample was used in training a model. The classical shadow-model formulation by Shokri et al. [5] established that models often exhibit measurable behavioral differences between member and non-member samples. Data-extraction research further demonstrates that LLMs can reveal portions of their training data through crafted prompts alone [14], indicating that observable traces of training data may exist.
However, recent large-scale evaluations of MIAs reinforce that traditional membership-based techniques face substantial challenges at LLM scale and often fail to deliver reliable signals in practical settings [15]. Additionally, MIAs accuracy degrades as dataset sizes grow, and becomes near random under IID conditions [8], [9].
II-B Canary-Based Memorization Analysis
Another line of work studies model memorization through canaries: synthetic or carefully chosen sequences inserted into training data to measure leakage [16]. Canary-based studies reveal how training dynamics influence memorization and have recently been revisited for LLMs [17]. But these techniques require controlled insertion of test sequences or access to probability distributions. Yet, many commercial LLMs are often developed with only web-based interfaces to output text-based responses (e.g. OpenAI’s GPT series [18], Google’s Gemini series [10], and Anthropic’s Claude series [11]).
II-C Dataset Inference
Beyond per-sample membership, Steinke et al. [19] propose one-run privacy auditing, aggregating statistical signals across many samples. Maini et al.  [9] propose the first formal dataset inference in LLMs, showing that it can be achieved by aggregating multiple MIAs output features and comparing their statistical behavior across datasets. Nevertheless, their methods depend on access to probabilities, loss, or perplexity, and require an auxiliary dataset drawn from the same distribution as the target set—assumptions that do not hold in practice.
III Threat Model and Problem Statement
@@
In this paper, we focus on detecting dataset misuse in LLM training. Following the settings of recent model ownership verification [20, 21], we assume a scheme with three participants involved, including $Accuser$, ArbiterArbiter, and $Responder$. AccuserAccuser owns a publicly-accessible dataset $\mathcal{D}$ and disallows using $\mathcal{D}$ to train LLMs for profits. $Responder$ owns a private LLM $\mathcal{M}$ and allows paid queries to $\mathcal{M}$. Believing that $Responder$ secretly fine-tuned $\mathcal{M}$ on $\mathcal{D}$, $Accuser$ makes an accusation of dataset misuse. $Arbiter$, a trusted third party, aims to verify the accusation.
@@
The problem of dataset inference is to determine if a given model $\mathcal{M}$ has been trained on a complete dataset $\mathcal{D}$. Inference in the case where $\mathcal{M}$ is trained on only a single sample from $\mathcal{D}$ does not constitute dataset inference and is considered out of scope. The formal definition was introduced in prior work [8], and we refer interested readers to that paper for further details.
IV Motivation and Key Observations
IV-A Previous Research
@@
Background: To achieve dataset inference, Maini et al. propose the first yet only method [9], which we name as DIAα for easy references. DIAα is built on top of MIAs [5] against LLMs [6, 7]. At the high level, MIA is a function $f_{\mathcal{M}}:\mathcal{X}\rightarrow\{0,1\}$. Given an LLM model $\mathcal{M}$ and an input $x$ from space $\mathcal{X}$, $f_{\mathcal{M}}$ determines if $x$ is in $\mathcal{M}$’s training set (i.e., member or not). A representative $f_{\mathcal{M}}$ is threshold-based: $f_{\mathcal{M}}=\mathbb{1}[\mathcal{S}(\mathcal{M},x)<\delta]$, where $\mathcal{S}$ is a score function like loss [22] and perplexity [14] and $\delta$ is a pre-determined threshold. There are also many other variants of $f_{\mathcal{M}}$, including lowest probability-based [6], perturbation based [8, 23], reference model based [9], and entropy based [14].
@@
DIAα Design: Observing that MIAs may present random accuracy when detecting membership of individual data samples (especially when the training set and testing set of $f_{\mathcal{M}}$ are IID, namely independent and identically distributed [24]), DIAα aggregates the results of multiple MIAs across different data samples to perform dataset inference. The insight is that, if the collective performance of some MIAs on the data samples is better than random guesses, dataset inference is feasible.
@@
Technically, to perform inference of a victim dataset $\mathcal{D}$ against a suspect model $\mathcal{M}$, DIAα gathers a validation dataset $\hat{\mathcal{D}}$ (which is IID with $\mathcal{D}$) and builds a set of diverse MIAs. Each MIA is trained with a training split from $\mathcal{D}$ as members and a training split from $\hat{\mathcal{D}}$ as non-members. Using the outputs of all MIAs as features and the membership status as labels, DIAα further trains a linear regression model, aiming to learn which MIAs are more helpful. Another split from $\mathcal{D}$, the testing split, is then fed to the linear model to obtain a membership likelihood value for each sample, forming a likelihood vector. Similarly, a likelihood vector can be obtained with the testing split from $\hat{\mathcal{D}}$. Given the two likelihood vectors, DIAα performs t-test [25] using the alternate hypothesis that $\mathcal{D}$ is used for training $\mathcal{M}$, or mathematically, the mean value of $\mathcal{D}$’s likelihood vector is higher than $\hat{\mathcal{D}}$’s.
@@
DIAα Restrictions: For successful dataset inference, DIAα requires MIAs to present a certain level of accuracy. To this end, DIAα assumes grey-box access to $\mathcal{M}$ for obtaining rich inputs for MIAs. Specifically, the MIAs adopted by DIAα all require intermediate outputs from $\mathcal{M}$ (probabilities, loss, perplexity, etc.) to predict a data sample’s membership (R1). This reduces DIAα’s practicability. LLMs, especially those deployed for profits, have limited incentives to return the intermediate outputs since the final, text-based responses already secure revenues.
@@
As DIAα needs to train a linear regressor to learn the relationship between member samples from the victim dataset and non-member samples from the validation dataset, it requires the availability of a validation dataset that is IID with the victim dataset (R2). This increases the burden of using DIAα. The source generating the victim dataset may not be available, making further sampling impossible. Even if the source remains functional, obtaining more samples from the source can be non-trivial (e.g., getting more books from the same author is not always possible). In order to compare DIAα with our proposed method later in the evaluation, we provide an additional IID validation dataset to satisfy its requirements. Under this relaxed condition, DIAα achieves a detection accuracy comparable to ours, as shown in Table VII.
IV-B Our Observations
@@
In principle, MIAs exploit group-level discrepancies between the training set and the non-training set. For instance, a representative threshold-based MIA [22] relies on a clear separation between the loss of training and non-training samples. When the group-level discrepancies do not emerge, MIAs fail. Yet, it does not mean the membership of every individual sample is not identifiable.
@@
TABLE I: Tainted samples identified from databricks-dolly-15k when used to fine-tune GPT-3.5-Turbo.
@@
Figure 1: A comparison of GPT-3.5-Turbo’s response before and after fine-tuning. The longest common sub-sequence shared with the oracle output is highlighted in green.
❶Study Setup: We fine-tune GPT-3.5-Turbo with databricks-dolly-15k [26], a dataset of 15,011 instruction-following records published by Databricks in April 2023. The training data of GPT-3.5-Turbo was up to September 2021 [27], ensuring databricks-dolly-15k was not included. We use the official APIs offered by OpenAI [28] to perform fine-tuning. We set the training epoch (n_epochs) to be 3, leaving all other parameters to default.
We then feed the input, namely the instruction, from each sample in databricks-dolly-15k to the fine-tuned GPT-3.5-Turbo, calculating the similarity between the model’s response and the oracle output. We adopt BERTScore [29] as the similarity metric, which leverages the pre-trained contextual embeddings from BERT [30] and matches words in the response and oracle ouput by cosine similarity. BERTScore offers two advantages. First, it has been shown to correlate with human judgment on sentence-level and system-level evaluation [31]. Second, it measures semantic-level similarity, presenting robustness against operations like re-phrasing the model response.
❷Results: In the study, we only consider samples resulting in a BERTScore over 0.95 to be tainted samples. Even using this conservative standard, we identify hundreds of tainted samples from the major categories in databricks-dolly-15k, as summarized in Table I. We also demonstrate an example in Figure 1, attaching the model’s responses before and after fine-tuning for references. Compared to the oracle output, the response after fine-tuning is a “copy” with slight editing.
@@
Figure 2: A databricks-dolly-15k example where GPT-3.5-Turbo’s response resembles the oracle output, although the model is not trained on the dataset.
TABLE II: “Fake” tainted samples identified from databricks-dolly-15k with non-fine-tuned GPT-3.5-Turbo.
@@
Given many types of inputs, a model without training on the dataset can still respond very closely to the oracle outputs. Figure 2 presents an illustrative example, which we may mistakenly classify as a tainted sample even using a high similarity threshold. Such cases are common. We redo the study presented in Observation I but using GPT-3.5-Turbo without fine-tuning. As summarized in Table II, hundreds of samples from databricks-dolly-15k are classified as tainted samples. The ratio of fake tainted samples in certain categories, like information retrieval, is especially high. In this case, the method above will wrongly report that GPT-3.5-Turbo used databricks-dolly-15k for training.
@@
Discussion: While we cannot reliably pinpoint individual tainted samples, we may still rely on the statistical results of tainted samples for dataset inference. As shown in Table I and Table II, the total number of tainted samples after fine-tuning is significantly higher than before fine-tuning. This inspires a new idea to determine if suspect model $\mathcal{M}$ used victim dataset $\mathcal{D}$ for training: if the amount of tainted samples identified with $\mathcal{M}$ is closer to the situation where $\mathcal{D}$ is involved for training, we report $\mathcal{M}$ used $\mathcal{D}$, and not otherwise.
V Our Methods
V-A Overview